OpenStudy (kydawg):

Verify the identity. cos(x+π/2) = -sin x

Vocaloid (vocaloid):

hint: cos(A+B) = cos(A)cos(B) - sin(A)sin(B) replace "A" with x and "B" with pi/2 and you'll see for yourself that it equals - sinx

OpenStudy (kydawg):

cos(x)cos(pi/2) - sin(x)sin(pi/2) = cosx(0) - sinx(1) = 0-sinx so -sinx? Did I do it right?!?!?!

Vocaloid (vocaloid):

very good!

OpenStudy (kydawg):

Thank you!